# Depth Restoration of Hand-Held Transparent Objects for Human-to-Robot Handover

### v1

@@ III-C H2R Handover Strategy of Transparent Objects
$C=w_{s}s+w_{r2h}\mathbf{v}_{r2h}^{T}(R_{g}\,\mathbf{v}_{r2h})+w_{u2d}\mathbf{v}_{u2d}^{T}(R_{g}\,\mathbf{v}_{u2d}),$ (4)
where $R_{g}$ is the rotation of grasp pose, $\mathbf{v}_{r2h}$ denotes the unit vector pointing from the robot to the human, and $\mathbf{v}_{u2d}$ denotes the unit vector pointing from the up to the down. We design the metric in this way because a grasp facing toward the user and vertically pointing downward is preferred. $w_{s}$, $w_{r2h}$, and $w_{u2d}$ are the weights, which are set to be $w_{s}=1$, $w_{r2h}=0.4$, and $w_{u2d}=0.2$. The grasp pose with the highest score will be chosen, and the corresponding pre-grasp pose is defined as 10cm back from the predicted grasp along its z-axis.
@@
We denote the initial and current hand keypoints as $X_{i}$ and $X_{c}$ in the dimension of $\mathbb{R}^{21\times 3}$. The translation vector is computed as $T=X_{c}^{0}-X_{i}^{0}$, where $X_{i}^{0}$ and $X_{c}^{0}$ represent the initial and current wrist position. For the rotation, we perform the Singular Value Decomposition on the covariance matrix $H$, defined in Equation 5,
$H=(X_{i}-X_{i}^{0})^{T}(X_{c}-X_{c}^{0})=U\Sigma V^{T}.$ (5)
The rotation matrix is then calculated as $R=VU^{T},$ where $R$ should be adjusted by flipping the sign of the last column if its determinant is negative. The translation and rotation are then applied to the initial grasp pose to obtain the current one.
Grasp & Retrieve. Once the robot reaches the pre-grasp position, it will start the grasp action in an open-loop manner. We move the final grasp position 5cm forward along the z-axis as we assume the human user will adapt to the robot’s grasp motion. Once the gripper closes successfully, the robot will retrieve and drop the object.
@@ IV-A Depth Restoration Experiments on TransHand-14K
Evaluation Metrics. We adopt four commonly used metrics for depth estimation [5]. All metrics are calculated on the transparent areas according to the given transparent masks.
  • RMSE: the root mean square error between the predicted result and ground truth.
  • REL: the mean value of absolute relative difference.
  • MAE: the mean absolute error between the predicted result and ground truth.
  • Threshold $\delta$: the percentage of data smaller than the given threshold $\delta$. $\max\left(d_{i}/d_{i}^{*},d_{i}^{*}/d_{i}\right)<\delta$, where $d$ represents the predicted depth, $d^{*}$ represents the ground truth depth, and $\delta\in\{1.05,1.10,1.25\}$.
Implementation Details. Image resolution is set to be 224 $\times$ 224 for both RGB and depth input. The RGB input is further augmented by adjusting the color space and incorporating additional elements such as bright patches, blur, and noise. We train the network for 100 epochs and choose Adam as the optimizer with a learning rate of 1e-3 for the first 80 epochs and 1e-4 for the last 20 epochs. The number of voxels used for restoration is $8^{3}$. We set $w_{\text{depth}}=200$, $w_{\text{prob}}=10$, and $w_{\text{norm}}=0.5$ in training.
Comparison with SOTA. We compare the proposed method with several state-of-the-art methods, including TODE-Trans [7], TransCG [16], SwinDRNet [8], and LIDF [15]. All baselines are trained on the proposed dataset TransHand-14K. We split the training, validation, and testing set according to the ratio of 7:2:1.
Table I reports the comparison results. We first compare methods on the known category objects, where models are trained and evaluated on all eight different shapes of transparent objects. Our method achieves the best performance in terms of RMSE, MAE, $\delta_{1.05}$, and $\delta_{1.10}$. For the other two metrics, our method is slightly weaker than SwinDRNet. For the unknown category evaluation, we choose six objects for training and the other two for evaluation. Our method achieves the best results in all metrics. It shows that by introducing hand pose information, the model’s generalization ability can be improved for unseen objects.
@@
Figure 7: Qualitative evaluation in real scenarios. We compare the proposed hand-ware depth model to state-of-the-art methods. We highlight defects including deformation (in the color of green) and spatial shifting (in the color of red). For better visualization, the point cloud of background is removed.
@@
TODE-Trans 0.024 0.056 0.017 57.03 82.55 98.61
TransCG 0.012 0.026 0.008 86.91 96.96 99.76
SwinDRNet 0.009 0.015 0.005 94.82 98.26 99.87
LIDF 0.014 0.029 0.010 85.18 96.08 99.60
@@
TODE-Trans 0.052 0.055 0.037 62.86 89.26 97.21
TransCG 0.027 0.064 0.020 51.54 78.54 98.79
SwinDRNet 0.022 0.049 0.015 65.07 86.79 98.79
LIDF 0.026 0.063 0.021 52.66 78.69 97.90
@@
TABLE II: Ablation Study: How hand-aware, multi-scale aggregation and point-wise feature fusion contribute to the results.
@@ IV-B Real-World Experiments
We conduct real-world experiments to verify the proposed transparent object depth restoration and handover method.
Experiment Setup. Fig. 6 shows the setting of our real robot system. A UR5 robot arm equipped with a soft gripper is rigidly mounted on the table, which is used to perform the robot handover task. For perception, we use a RealSense D435 camera with an RGB-D stream. The camera is placed in a third-person view with extrinsic calibrated. A human participant stands on the other side of the table, holding a transparent object in hand and passing it to the robot.
Implementation. In addition to implementing the proposed depth restoration method, some other perception models are used in the real-world experiment. To obtain the segmented area of transparent objects, we train a segmentation model in the Unet architecture with a ResNet-34 [29] as the encoder using the TransHand-14K dataset. For the hand pose detector, we find the weighted average regression from the depth image is vulnerable to occlusions and noise. Instead, we first generate 2.5D hand keypoints using an off-the-shelf RGB-based hand pose detector [31], and then extend it into 3D keypoints using the depth value of the wrist keypoint position. The pre-trained GraspNet-baseline [32] is used to generate 6-DoF grasping poses. All the aforementioned models are implemented using a computer with an NVIDIA 3080 GPU.
Qualitative Evaluation in Real Scenario. To better illustrate the sim-to-real performance of the proposed method, a qualitative evaluation is conducted by comparing it to the state-of-the-art methods in real scenarios. As shown in Fig. 7, it can be observed that the reconstructed object point cloud exists deformation (highlighted with green boxes) and spatial shifting (highlighted with red boxes). Results demonstrate our method can restore the depth value of transparent objects more accurately and robustly.
Handover Experiment. We conduct a human-to-robot handover experiment using the proposed handover pipeline. In the experiment, eight transparent objects with different shapes are used. To better illustrate the generalization ability of our model, three of them are in novel shapes. For each object, we perform six independent trials. The baseline chosen for comparison only uses GraspNet without depth restoration. In addition to assessing the overall success rate, we also compare the number of objects that exceed the specific success rate threshold $\delta\in\{0.5,0.8,1.0\}$.
@@
Table III reports the experimental results. Without depth restoration, GraspNet struggles to generate the correct grasp pose for handover due to the incomplete and inculcated depth value. The proposed depth restoration method can significantly improve the handover success rate, from 10.4% to 70.8%. Six among eight objects achieve a success rate of over 50%.
@@ V Conclusion
In this study, we introduce a novel solution for estimating the depth of hand-held transparent objects and highlight its significant impact on human-robot interaction. We present TransHand-14K, a synthetic dataset created through a real-to-sim data generation approach, effectively bridging the gap between simulated and real-world environments. Our proposed hand-aware model incorporates hand posture information to enhance depth restoration, demonstrating impressive generalization capabilities. Additionally, we have developed a handover pipeline that leverages our depth restoration technique. Evaluations of both the depth restoration process and real-world human-robot handover tasks validate the effectiveness of our approach.

### v2

@@ III-C H2R Handover Workflow of Transparent Objects
$C=w_{\text{s}}s+w_{r2h}\mathbf{v}_{\text{r2h}}^{\top}(R_{\text{g}}\,\mathbf{v}_{\text{r2h}})+w_{\text{u2d}}\mathbf{v}_{\text{u2d}}^{\top}(R_{\text{g}}\,\mathbf{v}_{\text{u2d}}),$ (4)
where $R_{\text{g}}$ is the rotation of grasp pose, $\mathbf{v}_{\text{r2h}}$ denotes the unit vector pointing from the robot to the human, and $\mathbf{v}_{\text{u2d}}$ denotes the unit vector pointing from the up to the down. We design the metric in this way because a grasp facing toward the user and vertically pointing downward is preferred. $w_{s}$, $w_{\text{r2h}}$, and $w_{\text{u2d}}$ are the weights, which are set to be $w_{\text{s}}=1$, $w_{\text{r2h}}=0.4$, and $w_{\text{u2d}}=0.2$. The grasp pose with the highest score will be chosen, and the corresponding pre-grasp pose is defined as 10cm back from the predicted grasp along its z-axis.
@@
We denote the initial and current hand keypoints as $X_{\text{i}}$ and $X_{\text{c}}$ in the dimension of $\mathbb{R}^{21\times 3}$. The translation vector is computed as $T=X_{\text{c}}^{0}-X_{\text{i}}^{0}$, where $X_{\text{i}}^{0}$ and $X_{c}^{0}$ represent the initial and current wrist position. For the rotation, we perform the Singular Value Decomposition on the covariance matrix $H$, defined in Equation 5,
$H=(X_{\text{i}}-X_{\text{i}}^{0})^{\top}(X_{\text{c}}-X_{\text{c}}^{0})=U\Sigma V^{\top}.$ (5)
The rotation matrix is then calculated as $R=VU^{\top},$ where $R$ should be adjusted by flipping the sign of the last column if its determinant is negative. The translation vector and rotation matrix are then applied to the initial pre-grasp pose to obtain the current one.
Grasp & Retrieve. Once the robot reaches the pre-grasp position, it will grasp the object in an open-loop manner. We move the final grasp position 5cm forward along the z-axis as we assume the human user will adapt to the robot’s grasp motion. Once the gripper closes successfully, the robot will retrieve and drop the object.
@@ IV-A Depth Restoration Experiments on TransHand-14K
Evaluation Metrics. We adopt four commonly used metrics for depth estimation [3]. 1) RMSE: the root mean square error between the predicted result and ground truth. 2) REL: the mean value of absolute relative difference. 3) MAE: the mean absolute error between the predicted result and ground truth. 4) Threshold $\delta$: the percentage of data smaller than the given threshold $\delta$. $\max\left(d_{i}/d_{i}^{*},d_{i}^{*}/d_{i}\right)<\delta$, where $d$ represents the predicted depth, $d^{*}$ represents the ground truth depth, and $\delta\in\{1.05,1.10,1.25\}$. All metrics are evaluated on the transparent objects area.
Implementation Details. We set the image resolution to be 224 $\times$ 224 for both RGB and depth input. The RGB input is further augmented by adjusting the color space and incorporating additional elements such as bright patches, blur, and noise. We train the network for 100 epochs and choose Adam as the optimizer with a learning rate of 1e-3 for the first 80 epochs and 1e-4 for the last 20 epochs. The number of voxels used for restoration is $8^{3}$. We set $w_{\text{depth}}=200$, $w_{\text{prob}}=10$, and $w_{\text{norm}}=0.5$ in training.
Comparison with SOTA. We compare the proposed HADR with several state-of-the-art methods, including TODE-Trans [5], TransCG [14], SwinDRNet [6], and LIDF [13]. All baselines are trained on the proposed dataset TransHand-14K. We split the training, validation, and testing set according to the ratio of 7:2:1.
Table I reports the comparison results. We first compare methods on the known category objects, where models are trained and evaluated on all eight different shapes of transparent objects. Our method achieves the best performance in terms of RMSE, MAE, $\delta_{1.05}$, and $\delta_{1.10}$. For the other two metrics, our method is slightly weaker than SwinDRNet. For the unknown category evaluation, we choose six objects for training and the other two for evaluation. Our method significantly outperforms other methods in terms of all metrics, demonstrating the strong generalization ability of the model. This better performance can be attributed to our hand-aware design, which enables the model to effectively transfer similar hand-object interaction patterns to unknown objects. Additionally, unlike LIDF [13], our method can directly produce high-quality final results without needing a depth refinement process.
@@
Figure 7: Qualitative evaluation in real scenarios. We compare the proposed HADR to state-of-the-art methods. Defects including deformation (in the color of green) and spatial shifting (in the color of red) are highlighted. For better visualization, we remove point cloud of the background.
@@
TODE-Trans [5] 0.024 0.056 0.017 57.03 82.55 98.61
TransCG [14] 0.012 0.026 0.008 86.91 96.96 99.76
SwinDRNet [6] 0.009 0.015 0.005 94.82 98.26 99.87
LIDF+Refine [13] 0.014 0.029 0.010 85.18 96.08 99.60
@@
TODE-Trans [5] 0.052 0.055 0.037 62.86 89.26 97.21
TransCG [14] 0.027 0.064 0.020 51.54 78.54 98.79
SwinDRNet [6] 0.022 0.049 0.015 65.07 86.79 98.79
LIDF+Refine [13] 0.026 0.063 0.021 52.66 78.69 97.90
@@
TABLE II: Ablation Study of Features and Modules.
@@ IV-B Real-World Experiments
We conduct real-world experiments to verify the effectiveness of HADR and handover workflow.
Experiment Setup. Fig. 6 illustrates the setting of our real robot system. A UR5 robot arm equipped with a soft gripper is rigidly mounted on the table, which is used to perform the robot handover task. For perception, we employ a RealSense D435 camera with an RGB-D stream. The camera is placed in a third-person view with extrinsic calibrated. A human participant stands on the opposite side of the table, holding a transparent object in hand and passing it to the robot. We choose the tablecloth with intricate patterns and rich colors as the background to examine the robustness of our method.
Implementation. In addition to implementing the proposed depth restoration method, some other perception models are used in the real-world experiment. To obtain the segmented area of transparent objects, we train a segmentation model in the Unet architecture with a ResNet-34 [27] as the encoder using the TransHand-14K dataset. For the hand pose detector, we find the weighted average regression from the depth image is vulnerable to occlusions and noise. Instead, we first generate 2.5D hand keypoints using an off-the-shelf RGB-based hand pose detector [29], and then extend it into 3D keypoints using the depth value of the wrist keypoint position. The pre-trained GraspNet-baseline [30] is used to generate 6-DoF grasping poses. All the aforementioned models are implemented using a computer with an NVIDIA 3080 GPU. In the experiment, the whole perception pipeline achieves an inference speed of 11 FPS, allowing real-time depth restoration.
Qualitative Evaluation in Real Scenarios. To better illustrate the sim-to-real performance of the proposed method, a qualitative evaluation is conducted by comparing it to the state-of-the-art methods in real scenarios. As shown in Fig. 7, it can be observed that the reconstructed object point cloud exists deformation (highlighted with green boxes) and spatial shifting (highlighted with red boxes). Results demonstrate that our method reconstructs transparent objects with greater accuracy and robustness in real-world settings.
Handover Experiment. We conduct a human-to-robot handover experiment using the proposed handover workflow. In the experiment, eight transparent objects with different shapes are used. To better illustrate the generalization ability of our model, three of them are in novel shapes. For each object, we perform six independent trials with different ways of holding. The baseline chosen for comparison is GraspNet without depth restoration. In addition to assessing the overall success rate, we also compare the number of objects that exceed the specific success rate threshold $\delta\in\{0.5,0.8,1.0\}$.
@@
Table III reports the experimental results. Without depth restoration, GraspNet struggles to generate the correct grasp pose for handover due to the incomplete and inaccurate depth value. The proposed depth restoration method can significantly improve the handover success rate, from 10.4% to 70.8%. Six among eight objects achieve a success rate of over 50%.
@@ V Conclusion
In this study, we introduce a novel solution for restoring the depth of hand-held transparent objects and highlight its great value in human-robot interaction. We present TransHand-14K, a synthetic dataset generated through a real-to-sim data generation scheme, effectively bridging the gap between simulated and real-world environments. Our proposed method, HADR, leverages hand posture information to enhance depth restoration, achieving superior reconstruction results and better generalization capabilities compared to existing methods. Additionally, we develop a handover workflow that leverages our depth restoration technique. Evaluations on both depth restoration and human-to-robot handover validate the effectiveness of our approach.